# Spatialyze: A Geospatial Video Analytics System with Spatial-Aware Optimizations

### v3

@@ 4.1.3. Movable Objects
Besides static properties like Geographic Constructs, a World may contain multiple Movable Objects. Being movable, they each have spatiotemporal properties, e.g., its location at a given time. They also have non-spatiotemporal properties including object ID and object type that does not change throughout the object’s life-span. For example, a Movable Object with object type of “car” will always be a “car.” Spatialyze models each Movable Object as $(oid,type,((l_{0},r_{0},p_{0},t_{0}),\allowbreak(l_{1},r_{1},p_{1},t_{1}),...))$, where $oid$ is its identifier; $type$ is the object type, such as car, truck, or human. At time $t_{i}$, $l_{i}=(x_{i},y_{i},z_{i})$ is its location; $r_{i}$ is its 3D rotation, defined as a quaternion (Kuipers, 1999); $p_{i}$ is its type-specific properties.
A Movable Object with $type=camera$ represents a camera that captures videos. It has one type-specific property (Zhang, 2014) $p_{i}=(it_{i},)$, where $it_{i}$ is the intrinsic (Anwar, 2022; Zhang, 2014) of the camera at time $t_{i}$. Despite being Movable Objects, this paper distinguishes camera from other types of Movable Objects with the following reason. Movable Objects with $type=camera$ are parts of users’ inputs into the World, while Spatialyze infers Movable Objects with other types from the visual contents captured by the cameras.
We do not include “video” as a part of our conceptual data model. A video itself does not have any geospatial property, so it does not exist in a World. Similarly, none of the objects tracked in the video exists initially in the World. Internally, only after we bind the video with a camera that captures the video, the video gains geospatial properties from the camera and become a part of the camera. Objects tracked in the video also gain their geospatial properties, as a result of the binding. Spatialyze derives the objects’ geospatial locations, by combining their 2 dimensional bounding boxes and camera’s geospatial information. An object that gains geospatial properties then becomes a Movable Object in the World.
@@ 4.2. S-Flow
We design S-Flow based on the data model presented in § 4.1, where users interact and manipulate geospatial objects via Geographic Constructs and Movable Objects, rather than video frames. S-Flow’s language constructs consists of functions that operate on World, RoadNetwork, and GeospatialVideo. Users compose such functions to construct geospatial video analytics workflows.
@@ 4.2.1. Camera
Lines LABEL:line:gsvideo-LABEL:line:end-build of Fig. 1 defines a Movable Object with $type=camera$ as discussed in in § 4.1.3. Users initialize a Camera from a list, where the $i$-th element of the list is a dictionary with 4 fields: translation, rotation, intrinsic, and timestamp. They correspond to $l_{i}$, $r_{i}$, $it_{i}$, and $t_{i}$, as defined in § 4.1.3.
@@ 4.2.2. GeospatialVideo
A GeospatialVideo is a video enhanced with geospatial metadata. A Video (lines LABEL:line:gsvideo-LABEL:line:end-build of Fig. 1) by itself contains a sequence of images where each represents a frame of the video. Such video exists spatially in the World after its Video object is bound to a Camera to obtain its spatial properties, where each frame $i$ is an image taken from the camera at location $l_{i}$, rotation $r_{i}$, intrinsic $it_{i}$, and at time $t_{i}$. The objects in each frame obtains its geospatial properties after its track is inferred from GeospatialVideo. These objects then become Movable Objects as defined in § 4.1.3.
@@ 4.2.3. RoadNetwork
A RoadNetwork is a collection of road segments. Each road segment is a Geographic Construct as defined in § 4.1.2. Besides being a Geographic Construct, a road segment stores its segment heading, which indicate traffic headings in that road segment. A road segment have multiple segment headings (curve lane) or no segment heading at all (intersection). Road segments in the RoadNetwork dataset that we use have following road types: roadsection, intersection, lane, and lanegroup. In S-Flow, we initialize a RoadNetwork by loading from a directory containing files of road network from each type. These road network information is accessible and often provided with AV video datasets (Fritsch et al., 2013; Caesar et al., 2019).
@@ 4.2.4. World
Build. A World is initially empty. Users add RoadNetwork using addGeogContructs(), following by adding GeospatialVideo via addVideo(). At this point of the workflow, the World contains Geographic Constructs, and Cameras. Spatialyze uses YOLOv5 (Jocher et al., 2022) and StrongSORT (Du et al., 2023) for object detection and object tracking by default. Spatialyze also supports using UDFs for object detection or tracking, as shown commented in line LABEL:line:world of Fig. 1.
Filter. Users then construct predicates to filter objects of their interest. As discussed in § 3, Spatialyze has not extracted detections and trackings of objects from the videos yet at this point. Nonetheless, Spatialyze provides an interface for users to interact with these objects as if they are already exist. For instance, object() refers to an arbitrary Movable Object with $type\neq camera$ in the World that is referenced in the predicate. Calling multiple object() gives users multiple arbitrary Movable Objects that appear together in the same video to use in the predicate. For existing information, camera() refers to an arbitrary Camera . Similarly, geoConstruct(type=...) refers to a Geographic Construct of a certain type. Users can construct a predicate involving multiple Movable Objects , a Camera, or Geographic Constructs, using our provided helper functions. For instance, contains(intersection,obj) determines if intersection as a polygon contains obj as a point. headingDiff(obj,cam,between=[0,9]) determines if the difference of the heading of obj and the heading of cam is between 0 and 9 degrees. In addition, we provide more helper functions and an interface for users to define their own helper functions. Finally, users can chain the filter or use boolean operators to combine predicates.
Observe. After filtering their desired Movable Objects, users observe the filtered World, using one of the following 2 observer functions. First, they can observe the World by watching through snippets of the users’ own input videos. Using saveVideos(file, addBoundingBoxes) function, Spatialyze saves all the video snippets that contains the Movable Objects of interest to a file, where addBoundingBoxes is Boolean flag that indicate whether the bounding box of each Movable Objects should be shown. Second, they can observe the World by getting the Movable Object directly. Using getObjects(), Spatialyze returns a list of Movable Objects.
@@ 5. Workflow Execution in Spatialyze
Figure 2. Geospatial Video Analytic Workflow Execution.
Figure 3. Processing Plan for the workflow shown in Fig. 1.
From users’ view, Spatialyze simply filters objects in the World after they call filter(). Then, they can observe the filtered objects afterwards, using one of the observer functions described in § 4.2.4. Internally, Spatialyze only performs video processing to detect objects and compute tracks for them as needed given the workflow. For example, in line LABEL:line:pred-heading of Fig. 1, the user filters on objects’ moving directions. Executing this requires computing the trajectory of objects in the given videos to compute their directions. Given an S-Flow workflow, the goal of Spatialyze is to understand the users’ workflow, recognize and execute the video processing operators that are needed for user’s filter predicates, and return the Movable Objects that satisfy users’ filters.
In § 4, we discuss a high-level interface for Spatialyze that allows users to declaratively construct their geospatial video analytic workflows. Users follow the build-filter-observe pattern without knowing when the actual execution happens. Internally, all execution happens when the users observe the world. In this section, we discuss the benefits of deferred execution and how Spatialyze generates execution plans for a workflow.
@@ 5.1. Challenges in Workflow Execution
We identify 3 challenges in building a geospatial video analytics system, specifically designing a system that is fast, capable of processing arbitrary-length videos, and extensible. First, processing videos with ML functions is slow; but Spatialyze can leverage its knowledge of the entire workflow in deciding which video frames and ML functions that must be executed. Second, one minute of a 1080p 12fps decoded video is 4.4 GB. Two days of such video can already fill up one GCP VM with the largest available memory size of 11,776 GB as of 2023. Spatialyze must scale with video sizes with limited memory. Finally, Spatialyze must be extensible with future ML functions and optimization techniques. We discuss how we address these challenges in § 5.2 and § 6.
@@ 5.2. System Design & Workflow Execution
To minimize execution time, Spatialyze defers ML inferences, which are expensive, until users observe the World. This way, Spatialyze can optimize the entire workflow execution. Once users observe a World, Spatialyze starts executing users’ workflow in 4 stages sequentially: (1) Data Integrator, (2) Video Processor, (3) Movable Objects Query Engine, and (4) Output Composer; as shown in Fig. 2.
@@ 5.2.2. Video Processor
This stage takes in users’ GeospatialVideo and users’ filter predicates from their workflow’s Filter. It then extracts Movable Objects from the GeospatialVideo. To do so, it first decodes the video to get image frames. It then executes an object detector to extract bounding boxes of objects from each frame. Based on the bounding box, image frame, and Camera, it estimates the 3D location of each object. At the end, it executes the object tracker on the bounding boxes and image frame to get Movable Objects. By default, Spatialyze uses an off-the-shelf algorithm to implement each step. The video decoder is implemented using OpenCV as it is the fastest option available. The object detector is implemented using YOLO (Jocher et al., 2022), a SOTA object detector used by our prior work (Romero et al., 2022; Xu et al., 2022) with high accuracy, low runtime, and capability to detect multiple types of objects. The object 3D location estimator is implemented using Monodepth2 (Godard et al., 2019) as it is the SOTA image depth estimation that only requires an image as an input. The object tracker is implemented using StrongSORT (Du et al., 2023) as it is a fast and online method, suited for processing large videos.
We designed the video processor to stream video frames through all the above algorithms. Each video frame is pipelined through decoding, object detection, 3D location estimation, and object tracking, hence our video processor only needs to keep $O(1)$ video frames and can handle videos of arbitrary size, addressing the second challenge. While stream processing of videos has been used previously (Broström, 2022), applying it to a streaming video processor is not trivial. For example, the original YOLO-StrongSORT object tracker (Broström, 2022) streams video frames. However, if we extended the tracker with our optimization techniques naively, we would have to hard-code them into the tracker’s implementation. This does not scale to future trackers. To solve this problem (and address the third challenge), we implemented our video processor as a plan of streaming operators shown in Fig. 4. Our baseline operators111 The streaming operators internally process videos and geospatial metadata to extract (2D or 3D) object detections and/or tracks. In contrast, users use predicate operators (e.g., contains, distance) to define filter predicates. include: (1) Video Decoder; (2) Object Detector; (3) 3D Location Estimator; and (4) Object Tracker. Specifically, (1) decodes each video to get a sequence of RGB images representing each frame of the video. (2) detects objects in each frame from (1), where each object and its object type and 2D bounding box are returned. AWorld is 3 dimensional where all objects have 3D locations; therefore, (3) estimates the 3D location of each object from (2) based on its Camera. Lastly, (4) tracks objects by associating detections from (3) between frames. We implemented each operator as an iterator function, where each operator takes in streams of per-frame inputs and return a stream of per-frame outputs. As a result, our video processor can process videos of arbitrary size, including those that are larger than the available memory. In addition, we implemented a caching system for the operators. In Fig. 4, frames is an input to both ObjectDetector and ObjectTracker. Instead of computing the stream of frames twice, Spatialyze iterates through frames once and caches the results. For instance, a frame $i$ is decoded when the ObjectDetector processes it. Spatialyze caches the frame $i$ so that the ObjectTracker can process it without having to decode the frame again. Spatialyze keeps track of the number of the operators that will be using a cached result and evicts a frame as soon as all of the operators finished processing it; therefore, Spatialyze only caches a small number of frames at a given time.

### v4

@@ 4.1.3. Movable Objects
Besides static properties like Geographic Constructs, a World may contain multiple Movable Objects. Being movable, each had spatiotemporal properties, e.g., its location at a given time. Each also has non-spatiotemporal properties, including object ID and object type, that do not change throughout the object’s life span. For example, a Movable Object with object type of “car” will always be a “car.” Spatialyze models each Movable Object as $(oid,type,((l_{0},r_{0},p_{0},t_{0}),\allowbreak(l_{1},r_{1},p_{1},t_{1}),...))$, where $oid$ is its identifier; $type$ is the object type, such as car, truck, or human. At time $t_{i}$, $l_{i}=(x_{i},y_{i},z_{i})$ is its location; $r_{i}$ is its 3D rotation, defined as a quaternion (Kuipers, 1999); $p_{i}$ is its type-specific properties.
A Movable Object with $type=camera$ represents a camera that captures videos. It has a type-specific property $p_{i}=(it_{i},)$, where $it_{i}$ is the intrinsic (Anwar, 2022; Zhang, 2014) of the camera at time $t_{i}$. Despite being Movable Objects, this paper distinguishes the camera from other types of Movable Objects for the following reason. Movable Objects with $type=camera$ are parts of users’ inputs into the World, while Spatialyze infers Movable Objects with other types from the visual contents captured by the cameras.
We do not include “video” as a part of our conceptual data model. A video itself does not have any geospatial property, so it does not exist in a World. Similarly, none of the objects tracked in the video exists initially in the World. Internally only after we bind the video with a camera that captures the video, the video gains geospatial properties from the camera and becomes a part of it. Objects tracked in the video also gain their geospatial properties, as a result of the binding, becoming a Movable Object in the World. Spatialyze derives the objects’ geospatial locations by combining their 2D bounding boxes and the camera’s geospatial information.
@@ 4.2. S-Flow
We design S-Flow based on the data model presented in § 4.1, where users interact and manipulate geospatial objects via Geographic Constructs and Movable Objects, rather than video frames. S-Flow’s language constructs consist of functions that operate on World, RoadNetwork, and GeospatialVideo. Users compose such functions to construct geospatial video analytics workflows.
@@ 4.2.1. Camera
Lines 3-3 of § 3 defines a Movable Object with $type=camera$ as discussed in in § 4.1.3. Users initialize a Camera from a list, where the $i$-th element of the list is a dictionary with 4 fields: translation, rotation, intrinsic, and timestamp. They correspond to $l_{i}$, $r_{i}$, $it_{i}$, and $t_{i}$, as defined in § 4.1.3.
@@ 4.2.2. GeospatialVideo
A GeospatialVideo is a video enhanced with geospatial metadata. A Video (lines 3-3 of § 3) by itself contains a sequence of images where each represents a frame of the video. Such video exists spatially in the World after its Video object is bound to a Camera to obtain its spatial properties, where each frame $i$ is an image taken from the camera at location $l_{i}$, rotation $r_{i}$, intrinsic $it_{i}$, and at time $t_{i}$. An object in each frame obtains its geospatial properties after its track is inferred from GeospatialVideo. These objects then become Movable Objects as defined in § 4.1.3.
@@ 4.2.3. RoadNetwork
A RoadNetwork is a collection of road segments. Each road segment is a Geographic Construct as defined in § 4.1.2. Besides being a Geographic Construct, a road segment stores its segment heading, which indicates traffic headings in that road segment. A road segment can have multiple segment headings (curve lane) or no segment heading (intersection). Road segments in the RoadNetwork dataset that we use have the following road types: roadsection, intersection, lane, and lanegroup. In S-Flow, we initialize a RoadNetwork by loading from a directory containing files of the road network from each type. This road network information is accessible and often provided with AV video datasets (Fritsch et al., 2013; Caesar et al., 2019).
@@ 4.2.4. World
Build. A World is initially empty. Users add RoadNetwork using addGeogContructs(), followed by adding GeospatialVideo via addVideo(). At this point of the workflow, the World contains Geographic Constructs, and Cameras. Spatialyze uses YOLOv5 (Jocher et al., 2022) and StrongSORT (Du et al., 2023) for object detection and object tracking by default. Spatialyze also supports using UDFs for object detection or tracking, shown as the comment in line 3 of § 3.
Filter. Users then construct predicates to filter objects of their interest. As discussed in § 3, Spatialyze has not extracted detections and trackings of objects from the videos yet at this point. Nonetheless, Spatialyze provides an interface for users to interact with these objects as if they already exist. For instance, object() refers to an arbitrary Movable Object with $type\neq camera$ in the World to use in the predicate. Calling multiple object() gives users multiple arbitrary Movable Objects that appear together in the same video. For existing information, camera() refers to an arbitrary Camera. Similarly, geoConstruct(type=...) refers to a Geographic Construct of a certain type. Users can construct a predicate involving multiple Movable Objects, a Camera, or Geographic Constructs, using our provided helper functions. For instance, contains(intersection,obj) determines if intersection as a polygon contains obj as a point. headingDiff(obj,cam,between=[0,9]) determines if the difference of the heading of obj and the heading of cam is between 0 and 9 degrees. In addition, we provide more helper functions and an interface for users to define their own helper functions. Finally, users can chain the filter or use boolean operators to combine predicates.
Observe. After filtering their desired Movable Objects, users observe the filtered World, using one of the following 2 observer functions. First, they can observe the World by watching through snippets of the users’ own input videos. Using saveVideos(file,addBoundingBoxes) function, Spatialyze saves all the video snippets that contain the Movable Objects of interest to a file, where addBoundingBoxes is a Boolean flag that indicates whether the bounding box of each Movable Objects should be shown. Second, they can observe the World by getting the Movable Objects directly. Using getObjects(), Spatialyze returns a list of Movable Objects.
@@ 5. Workflow Execution in Spatialyze
[…]
Spatialyze System Image: geospatial video analytic workflow execution. The user writes an S-Flow workflow in Python along with the input videos, cameras, and road network. Spatialyze’s Data Integrator integrates the videos and cameras into geospatial videos and ingests the road network into the geospatial metadata store. Its Video Processor processes the geospatial videos to get the Movable Objects. Its Movable Objects Query Engine filters the relevant Movable Objects based on the user’s filter. Finally, its Output Composer composes either the videos with the relevant Movable Objects or Object Trajectories for the user.
Spatialyze optimizes a video processing plan by inserting optimization operators into the plan. The Road Visibility Pruner is before the Video Decoder. The Object Type Pruner is right after the Object Detector. The Geometry-Based 3D Location Estimator replaces the ML-based 3D Location Estimator. The Exit Frame Sampler is right before the Object Tracker.
[…]
From users’ view, Spatialyze filters objects in the World after they call filter(). Then, they can observe the filtered objects afterward, using one of the observer functions described in § 4.2.4. Internally, Spatialyze only performs video processing to detect objects and compute tracks for them as needed given the workflow. For example, in line 3 of § 3, the user filters on objects’ moving directions. Executing this requires computing the trajectory of objects in the videos to compute their directions. Given an S-Flow workflow, the goal of Spatialyze is to understand the users’ workflow, recognize and execute only the video processing operators needed for users’ filter predicates, and return Movable Objects that satisfy the filters.
In § 4, we discuss a high-level interface for Spatialyze that allows users to construct their geospatial video analytics workflows declaratively. Users follow the build-filter-observe pattern without knowing when the actual execution happens. Internally, all execution happens when the users observe the world. In this section, we discuss the challenges in designing Spatialyze system and how Spatialyze generates execution plans for a workflow.
@@ 5.1. Challenges in Workflow Execution
We identify 3 challenges in building a geospatial video analytics system, specifically designing a system that is fast, capable of processing arbitrary-length videos, and extensible. First, processing videos with ML functions is slow; Spatialyze must leverage its knowledge of the entire workflow to decide which video frames and ML functions must be executed. Second, one minute of a 1080p 12fps decoded video is 4.4 GB. Two days of such video can already fill up one GCP VM with the largest available memory size of 11,776 GB as of 2023. Spatialyze must scale with video sizes with limited memory. Finally, Spatialyze must be extensible with future ML functions and optimization techniques. We discuss how we address these challenges in § 5.2 and § 6.
@@ 5.2. System Design & Workflow Execution
[…]
@@ 5.2.2. Video Processor
This stage takes in users’ GeospatialVideo and users’ filter predicates from their workflow’s Filter. It then extracts Movable Objects from the GeospatialVideo. To do so, it first decodes the video to get image frames. It then executes an object detector to extract bounding boxes of objects from each frame. Based on the bounding box, image frame, and Camera, it estimates the 3D location of each object. Finally, it executes the object tracker on the bounding boxes and image frame to get Movable Objects. By default, Spatialyze uses an off-the-shelf algorithm to implement each step. The video decoder is implemented using OpenCV (Bradski, 2000) as it is the fastest option available. The object detector is implemented using YOLO (Jocher et al., 2022), a SOTA object detector used by our prior work (Romero et al., 2022; Xu et al., 2022) with high accuracy, low runtime, and capability to detect multiple types of objects. The object 3D location estimator is implemented using Monodepth2 (Godard et al., 2019) as it is the SOTA image depth estimation that only requires an image as an input. The object tracker is implemented using StrongSORT (Du et al., 2023) as it is a fast and online method, suited for processing large videos.
[…]
@@
{listing}
[h!]{minted}[linenos,mathescape,escapeinside=!!,fontsize=]python frames = VideoDecoder(’videofile.mp4’) !! object2Ds = ObjectDetector(frames) !! object2DAnd3Ds = Loc3DEstm(frames, object2Ds, cameras) !! movableObjects = ObjectTracker(object2DAnd3Ds, frames) !! Baseline Video Processing Plan
Executing the plan computes 3D tracks of objects in a video. However, streaming each video through all operators is computationally expensive and not always necessary. Based on the users’ filter predicates from when they Filter the World, Spatialyze only includes the necessary operators in the execution plan. For instance, a predicate based on object types like obj.type=="car" requires objects’ types and hence the execution plan will include lines 5.2.2 and 5.2.2. Likewise, a predicate with distance or contain functions requires objects’ 3D location and therefore include lines 5.2.2, 5.2.2, and 5.2.2 in the plan. Finally, a predicate with a headingDiff function requires objects’ moving directions and include all lines in the plan.
We implemented each optimization technique in § 6 as a streaming operator. Two of the optimization operators execute users’ predicate early on to reduce the number of inputs (video frames and objects) into later operators. We designed our video processor that constructs plans using streaming operators; therefore, future optimization techniques and video processing functions can be added, addressing the runtime efficiency and extensibility challenges.
5.2.3. Movable Objects Query Engine
This stage takes in users’ filter predicates from their workflow’s Filter and resulting Movable Objects from the video processor. It then filters the Movable Objects with the users’ predicates in 2 steps. First, it streams the Movable Objects into a newly created table in the geospatial metadata store. The table has a spatial index on objects’ tracks and a temporal index on the period, in which the objects exist. Then, it translates the user’s filter predicates into an SQL query and executes it in the geospatial metadata store against the Movable Objects, Cameras, and RoadNetwork. When the user’s predicates involve multiple objects (e.g., distance(car1,car2)<5), Spatialyze self-joins the Movable Objects table to find all objects that coexist in a given period using the temporal index to speed up the join. When the predicates involve spatial relationships between a Movable Object and a RoadNetwork (e.g., contains(road,car)), the spatial index speeds up the joins between the Movable Objects and RoadNetwork tables. We implemented the metadata store using MobilityDB (Zimányi et al., 2020) as it provides necessary spatial and temporal indexes for movement objects. The index along with its spatiotemporal data type are suited for processing queries on Movable Objects and Geographic Constructs.
5.2.4. Output Composer
This stage composes the Movable Objects query results to users’ preferred formats for them to observe. Calling getObjects, Spatialyze returns a list of Movable Objects. Calling saveVideos, Spatialyze encodes the frames containing the Movable Objects into video files with colored bounding boxes annotation.
Fig. 1 shows the execution stages for the workflow in § 3. It contains all the stages mentioned above. A user first builds the world by adding data. Spatialyze’s Data Integrator creates an indexed table consisting of RoadNetwork and joins the Video and its corresponding Camera. Then, Spatialyze’s Video Processor creates an optimized video processing plan with 4 processing operators and 4 optimization operators (to be discussed in § 6). Parts of the users’ predicates are executed in Road Visibility Pruner and Object Type Pruner. The Movable Objects Query Engine streams the Movable Objects from the Video Processor into the geospatial metadata store and filters them with the users’ predicates. Finally, as the workflow asks for videos, Spatialyze formats and saves the frames that have the filtered objects into video snippet files.
6. Video Processing Optimization
SOTA video inference uses ML models to extract the desired information, such as objects’ positions and trajectories within each video. Running such models is computationally expensive. Spatialyze’s video processor operates with 4 optimization techniques to reduce such runtime bottlenecks. To be discussed below, the idea behind our optimization is to leverage the geospatial properties of the videos, and the Movable Objects’ inherited physical behaviors. We implement each optimization as a streaming operator in the Video Processing stage in § 5.2.2. In addition to the existing 4 processing operators, Spatialyze may place optimization operators in the middle of the video processing plan to reduce the amount of input into later processing operators based on the filter predicates. In § 6.1, Spatialyze places the Road Visibility Pruner after the Video Decoder, when users specify the Geographic Constructs the objects of interest must be in. Doing so reduces the number of input frames for the rest of the operators in the plan. In § 6.2, If users filter on object types, then Spatialyze places the Object Type Pruner right after the Object Detector. It reduces the number of objects that Spatialyze needs to calculate for their 3D locations and tracking. In § 6.3, Spatialyze replaces the ML-based 3D Location Estimation operator with the much faster 0pt, if Spatialyze can assume that the objects of interest are on the ground, such as cars or people. In § 6.4, Spatialyze places the Exit Frame Sampler in between the 3D Location Estimator and the Object Tracker, if the users’ predicates focus on cars. It utilizes the 3D locations of the objects and the user-provided metadata to reduce the number of frames that the Object Tracker processes.
Fig. 1 shows where the video processor executes all the optimization steps. Line 3 in § 3 corresponds to Road Visibility Pruner. Line 3 filters by object type, so Spatialyze adds Object Type Pruner. The objects of interest are cars or trucks, so Spatialyze adds Exit Frame Sampler and 0pt. In the following sections, we discuss each optimization in detail.
6.1. Road Visibility Pruner
In our experiments, the runtime of ML models (Jocher et al., 2022; Zhou et al., 2019; Du et al., 2023) in the video processing plan takes 90% of the entire workflow execution runtime, on average. Hence, the goal of the Road Visibility Pruner is to remove frames that do not contain objects of the user’s interest to reduce the overall runtime of the video processing plan.
6.1.1. High-Level Concepts
The Road Visibility Pruner uses Geographic Construct’s visibility as a proxy for a Movable Object’s visibility. Specifically, a predicate includes contains(road,obj) and distance(cam,obj)<d means that obj is not visible in the frames where road is not visible within d meters; therefore, Spatialyze prunes out those frames. In § 3, the Road Visibility Pruner partially executes lines 3 where the user searches for “cars within 50 meters at an intersection.” Spatialyze removes all frames that do not contain any visible intersection within 50 meters as these frames will not contain any car of users’ interest.
The Road Visibility Pruner only concerns the visibility of Geographic Constructs. Therefore, it uses Geographic Constructs and Camera to decide whether to prune out a video frame. It uses the intrinsic and extrinsic properties of Camera (Zhang, 2014) to determine the viewable area of each of the cameras according to the world coordinate system. [^2] World Coordinate System is a 3-dimensional Euclidean space, representing a World. It then determines if the viewable area of each frame includes any Geographic Construct of interest; frames that do not contain such constructs are dropped and will not be processed in the later processing operators. Since the Road Visibility Pruner will be the first processing operator if applied, it can reduce the number of frames that need to be processed by subsequent ML models in the plan, and hence can substantially reduce overall execution time.
6.1.2. Algorithm
The Road Visibility Pruner works on a per-frame basis. For each video frame $i$ shot by a Camera at location $l_{i}$ with rotation $r_{i}$ and camera intrinsic $it_{i}$, the Road Visibility Pruner performs three steps to determine if the frame should be pruned out. First, it computes the 3D viewable space of the Camera at frame $i$. Then, it uses the viewable space to find Geographic Constructs that are visible to the Camera at frame $i$. Finally, it uses the users’ filter predicates to determine if any Geographic Construct of interest is visible in the frame. We explain each step below.